OpenStudy (peterg2001):

will medal The following table shows a direct variation. 3 15 18 27 y 5 6 9 Find y. 0 1 2 3

OpenStudy (peterg2001):

@lizz123

OpenStudy (lizz123):

27/3=9 18/3=6 15/3 = 5 What does 3 divided by 3 equal

OpenStudy (lizz123):

\[3\div3\]

OpenStudy (lizz123):

It equals 1 y=1

OpenStudy (lizz123):

@peterg2001